KTeV Shift Summary for DAY 9/26/99 (Sun.)

Earl C. Swallow (earls@elmhurst.edu)
Sun, 26 Sep 1999 15:33:56 -0500 (CDT)


Shift Summary for DAY shift, 09/26/99 (Sun.)

On Shift: Ed Blucher & Earl Swallow + Arun, BobT, JasonH,
Ashkan & Theo(from Madison), Sydney, DougJ....

Hours of beam available at all intensities: 7.3 (out of 8.5)
Hours of beam available at nominal intensity: 7.3
Hours of E799 data (special runs not included): 6.8
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time & intensity:

Fairly frequent aborts, energy conservation periods, NM2EU current
limit work, Run end/start activities,

========================================================================
Runs ended this shift:

14735 -- BEAM_799 -- 16.8M
(with 2x prescale on Trigger 10)

Tapes brought to FCC this shift:

UP(A,B,C)025-027, UPS009 (runs through 14713-14734)
UP(A,B,C)028-030, UPS010 (Run 14735)

========================================================================
Accesses during shift:

None.

========================================================================
Overview of shift:

E799 data taking at 9E12 with 10" Be common absorber.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14735 07:00 Shift Begins. Run 14735 is in progress.

07:30 ECS changes the EPICURE window to "Reverse Video"
mode to get conventional EPICURE window "look" and
starts second window for running RONALD Scripts.
(Wish I knew how I got Launcher to come up so I
could start new window.)

07:40 Main SDAQ process is producing an error which causes
an automatic SDAQ restart -- sent note to Theo &
Ashkan.

08:20 MCR calls to say the low current limit on the
(recently replaced) NM2EU power supply is causing
aborts -- controls tech coming in to fix it

10:02 Beam off to work on NM2EU power supply current
limit.

10:30 Ashkan calls - he noticed that DC T0s for last
several runs have been shifted by ~ 1 nsec. He asks
that we install new T0 constants at the end of this
run.

10:50 Beam back at full intensity -- in fact nearly 1E13.

~230 DC4X trips -- reset by spill 234.
It tool a while to figure out why we had a
voiceless "beep."

11:00 CsI 1566 has +2.5% pedestal shift.

11:20 Arun deleted Online Split files which have already
been spooled to tape to prevent a disk space crisis.

~12:00 Scaler display totally freezes. Talk to Ashkan and
he restarts process by hand. It now works for a
few pulses and then the SDAQ system has to restart
it....

12:25 Mounted UPK003 for online split.

12:40 We also note that the SDAQ occasionally indicates an
RF error condition.

14:00 CPU usage for monitoring plane ~75-76%/spill. This
may be limitation at high intensity rather than
memory. Increased B02, B08 monitor prescales from 8
to 10.

14:20 CsI 2343 consistently has wide laser rms -- it
has 231 bad capids for the run.

14:30 Plot on page 296 of KTeV Logbook XXI claims first
observation of KL ---> e e mu mu from KTeV99.

14735 14:46 Taku calls Run 14735 to end.

14:50 Install new CSQ6 and DCT0 constants.

14:51 Plot of KL ---> pi pi e e events from KTeV99 appears
on page 297 of Logbook XXI (from MikeA).

15:00 NM2SR still does NOT reverse through EPICURE
(contrary to claim made to 9/25/99 EVE shift).

14736 15:05 Start Run 14736.

15:30 Shift ends with capid problems under investigation
with attendant mysteries. Stay tuned for the next
episode (EVE).

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Monday, DAY : PLEASE:

(1) put the old HP4Si toner cartridges (at least 4 in control room) into
the FNAL recycling system.

(2) get the "BAD" pieces of electronics which litter the area moved over
to PREP for service.

--
===============================================================================
Earl C. Swallow                           email:  earls@elmhurst.edu
Theophil W. Mueller Professor              (or)   earls@fnal.gov
   and Chairperson
Department of Physics                     Phone:  (630)-617-3577
Elmhurst College                          FAX:    (630)-617-3735
190 Prospect Avenue
Elmhurst, IL   60126-3296                 http://www.elmhurst.edu/~earls
===============================================================================
"Books are lighthouses erected in the dark sea of time."
                                --  from Gargoyles  (The Scrolls of Merlin)